# Fin-R1: A Large Language Model for Financial Reasoning through Reinforcement Learning

### v3

@@ 4.1 Supervised Fine-tuning
By definition, the use of the cross-entropy loss function (4.1) encourages the model to assign higher probability to the high-quality answers provided in the dataset. In doing so, it guides the model to internalize correct reasoning and answer patterns, effectively adapting it to the target domain task. In practice, the minimization of (4.1) is carried out via stochastic gradient optimization where gradients are obtained through backpropagation, and the parameters are updated using adaptive algorithms such as Adam (kingma2014adam).
@@ 4.2 Group Relative Policy Optimization
To compute the accuracy reward, we employ Qwen2.5-Max (qwen2024) as the evaluator. Specifically, for each candidate output $o$, we extract the answer enclosed within the $\ldots</answer>$ tags and prompt Qwen2.5-Max to assess whether the enclosed answer is semantically consistent with the ground-truth answer $s^{*}$ (See Section B of the Supplementary Material for the prompts used to instruct Qwen2.5-Max as the evaluator). If they are consistent, the reward is set to $1$; otherwise, it is set to 0. Formally, we define
@@
Advantage Calculation. Based on the reward $r_{i}$, PPO estimates the value function $V^{\pi_{\theta}}(x)=\mathbb{E}(r_{i}\mid x)$, where the expectation is taken over the stochasticity in generating the output induced by the policy $\pi_{\theta}$, which is typically random. It then calculates the advantage function as the difference between $r_{i}$ and the estimated value to construct the gradient to update $\theta$. The purpose of using the advantage rather than the raw reward is to reduce the variance of the gradient estimate, which in turn improves the statistical efficiency of policy learning (williams1992simple). However, the value function is typically parameterized by a transformer network, whose training and storage incur substantial computational and memory overhead.
@@ 5.1 Evaluation Datasets and Methodology
For numerical calculation questions, LLM-generated answers may be mathematically correct but expressed in a different valid format, due to variations in decimal precision or alternative numeric representations (see Figure 8 for illustrations). Directly comparing these answers with the reference answers will often mark them as incorrect because they do not match exactly, even though they should be considered correct. To address this, we use LLMs as an automated evaluation judge for answer check by adopting the prompt template and the evaluation methodology proposed by zhu2024judgelm. Although this evaluation paradigm appears straightforward, we have systematically conducted numerical experiments to fine-tune several critical parameters in the prompt template for optimizing the reliability of the LLM-based judge. Further experimental details can be found in the Section B of the Supplementary Material. Based on the optimized prompt, we instruct the LLM to report each model’s performance on a 100-point scale, which reflects the percentage of questions answered correctly by each model.
@@ 5.2 Baselines
We compare our Fin-R1 against eight state-of-the-art LLMs, including (i) DeepSeek-R1; (ii) DeepSeek-R1-Distill-Qwen-7B; (iii) DeepSeek-R1-Distill-Qwen-14B; (iv) DeepSeek-R1-Distill-Qwen-32B; (v) DeepSeek-R1-Distill-Llama-70B; (vi) Qwen-2.5-7B-Instruct; (vii) Qwen-2.5-14B-Instruct; and (viii) Qwen-2.5-32B-Instruct. The models labeled “Distill” are lightweight variants obtained through knowledge distillation, a technique that transfers knowledge from a large, high-performance teacher model (e.g., DeepSeek-R1) to smaller student models (e.g., Qwen and Llama series), while preserving comparable reasoning capabilities (see, e.g., Section 2.4 in guo2025deepseek). This approach enables the efficient deployment of advanced financial LLMs in resource-constrained settings. These baseline models were chosen to cover a broad range of sizes, from lightweight to large-scale architectures, while maintaining strong performance in terms of reasoning capability and computational efficiency. Our comparison aims to comprehensively evaluate Fin-R1 against leading baselines in financial applications.
@@ 6 Conclusion
THIS SUPPLEMENT IS STRUCTURED as follows. Section A systematically compiles information on the sources of the datasets used in our experiments and case studies, covering dataset categories, content descriptions, data sources, open-source status, language types, data types, and the sizes of the original and processed data; it also provides a structured table summarizing the key datasets employed in our financial AI experiments. Section B presents prompt templates for data distillation and filtering: Subsection B.1 introduces the data-distillation prompt, which follows the official DeepSeek-R1 configuration and comprises the task description, task input, task instruction, an explicit execution directive, and normative notes, together with the corresponding figure; Subsections B.2 and  B.3 outline the answer-check prompt and the reasoning-selection prompt, respectively. The former consists of a role definition, an input module, judgment rules for consistency assessment, and output-format requirements; the latter is designed around seven key dimensions and specifies evaluation criteria, input components, and the scoring method. Both prompts are accompanied by supporting figures. Subsection C.1 evaluates the impact of five prompt formats on evaluation-model performance, builds a quantitative evaluation metric system, and accordingly identifies Qwen2.5-72B-Instruct as the judge model for answer check; the experimental data and performance-comparison results both support this conclusion. Subsection C.2 reports supplementary experiments conducted to select the reasoning scoring model: we compare the scoring outcomes of Qwen2.5-72B-Instruct, GPT-4o, and human annotators, and visualize correlations among their score distributions using heatmaps. Section D presents a case study that intuitively demonstrates Fin-R1’s superior reasoning and response performance in financial scenarios.
@@ Appendix A Details for raw datasets
  • FinanceQT (malik2024quanttradingxu2023): Focuses on generating financial code and quantitative strategy scripts for financial scenarios.
  • Finance-500K (flowers2025financeinstruct): Covers professional financial content such as financial terminology explanation and QA on financial expertise.
  • FinanceIQ (financeIQ2023): Used for tasks including financial terminology explanation, QA on financial expertise, and financial calculations.
@@
  • Ant-Finance (alipay2023financial): Involves financial business tasks that do not require complex reasoning, such as content generation and compliance management.
  • FinCorpus (duxiaoman2023fincorpus): Contains corpora related to financial domain knowledge acquisition and financial text generation.
  • FinQA (chen-etal-2021-finqaa): Focuses on reasoning tasks for financial data, such as numerical reasoning in financial business.
  • ConvFinQA (chen2022convfinqa): Used for QA tasks in the financial field that require context-aware reasoning.
  • TFNS (anonymous2024twitter): Mainly supports reasoning tasks with sentiment analysis attributes, such as sentiment classification of financial news.
  • FinCUGE (lu2023bbtfin): Involves multiple types of reasoning tasks in the financial field, including numerical analysis and causal relationship extraction.
@@ B.3 The Prompt of Reasoning Selection
For reasoning selection in data filtering step, the prompt template presented in Figure 12 is designed based $7$ key dimensions (xie2024finnlpa) and is consisted of the evaluation criteria that structure how the reasoning process should be assessed; the inputs, the reasoning process and the standard answer, against which the evaluation is carried out; the description of scoring method, where each criterion is worth one point and the final decision is binary (1 for high-quality reasoning if all seven points are met, otherwise 0), with the score required for standardized extraction.

### v4

@@ 4.1 Supervised Fine-tuning
By definition, the use of the cross-entropy loss function (4.1) encourages the model to assign higher probability to the high-quality answers provided in the dataset. In doing so, it guides the model to internalize correct reasoning and answer patterns, effectively adapting it to the target domain task. In practice, the minimization of (4.1) is carried out via stochastic gradient optimization where gradients are obtained through backpropagation, and the parameters are updated using adaptive algorithms such as Adam (Kingma and Ba, 2014).
@@ 4.2 Group Relative Policy Optimization
To compute the accuracy reward, we employ Qwen2.5-Max (Team, 2024b) as the evaluator. Specifically, for each candidate output $o$, we extract the answer enclosed within the $\ldots</answer>$ tags and prompt Qwen2.5-Max to assess whether the enclosed answer is semantically consistent with the ground-truth answer $s^{*}$ (See Section B of the Supplementary Material for the prompts used to instruct Qwen2.5-Max as the evaluator). If they are consistent, the reward is set to $1$; otherwise, it is set to 0. Formally, we define
@@
Advantage Calculation. Based on the reward $r_{i}$, PPO estimates the value function $V^{\pi_{\theta}}(x)=\mathbb{E}(r_{i}\mid x)$, where the expectation is taken over the stochasticity in generating the output induced by the policy $\pi_{\theta}$, which is typically random. It then calculates the advantage function as the difference between $r_{i}$ and the estimated value to construct the gradient to update $\theta$. The purpose of using the advantage rather than the raw reward is to reduce the variance of the gradient estimate, which in turn improves the statistical efficiency of policy learning (Williams, 1992). However, the value function is typically parameterized by a transformer network, whose training and storage incur substantial computational and memory overhead.
@@ 5.1 Evaluation Datasets and Methodology
For numerical calculation questions, LLM-generated answers may be mathematically correct but expressed in a different valid format, due to variations in decimal precision or alternative numeric representations (see Figure 8 for illustrations). Directly comparing these answers with the reference answers will often mark them as incorrect because they do not match exactly, even though they should be considered correct. To address this, we use LLMs as an automated evaluation judge for answer check by adopting the prompt template and the evaluation methodology proposed by Zhu et al. (2024). Although this evaluation paradigm appears straightforward, we have systematically conducted numerical experiments to fine-tune several critical parameters in the prompt template for optimizing the reliability of the LLM-based judge. Further experimental details can be found in the Section B of the Supplementary Material. Based on the optimized prompt, we instruct the LLM to report each model’s performance on a 100-point scale, which reflects the percentage of questions answered correctly by each model.
@@ 5.2 Baselines
We compare our Fin-R1 against eight state-of-the-art LLMs, including (i) DeepSeek-R1; (ii) DeepSeek-R1-Distill-Qwen-7B; (iii) DeepSeek-R1-Distill-Qwen-14B; (iv) DeepSeek-R1-Distill-Qwen-32B; (v) DeepSeek-R1-Distill-Llama-70B; (vi) Qwen-2.5-7B-Instruct; (vii) Qwen-2.5-14B-Instruct; and (viii) Qwen-2.5-32B-Instruct. The models labeled “Distill” are lightweight variants obtained through knowledge distillation, a technique that transfers knowledge from a large, high-performance teacher model (e.g., DeepSeek-R1) to smaller student models (e.g., Qwen and Llama series), while preserving comparable reasoning capabilities (see, e.g., Section 2.4 in Guo et al., 2025a). This approach enables the efficient deployment of advanced financial LLMs in resource-constrained settings. These baseline models were chosen to cover a broad range of sizes, from lightweight to large-scale architectures, while maintaining strong performance in terms of reasoning capability and computational efficiency. Our comparison aims to comprehensively evaluate Fin-R1 against leading baselines in financial applications.
@@ Appendix A Details for raw datasets
  • FinanceQT (Malik, 2024): Focuses on generating financial code and quantitative strategy scripts for financial scenarios.
  • Finance-500K (Flowers, 2025): Covers professional financial content such as financial terminology explanation and QA on financial expertise.
  • FinanceIQ (Team, 2023b): Used for tasks including financial terminology explanation, QA on financial expertise, and financial calculations.
@@
  • Ant-Finance (Team, 2023a): Involves financial business tasks that do not require complex reasoning, such as content generation and compliance management.
  • FinCorpus (Team, 2023c): Contains corpora related to financial domain knowledge acquisition and financial text generation.
  • FinQA (Chen et al., 2021): Focuses on reasoning tasks for financial data, such as numerical reasoning in financial business.
  • ConvFinQA (Chen et al., 2022): Used for QA tasks in the financial field that require context-aware reasoning.
  • TFNS (Anonymous, 2024): Mainly supports reasoning tasks with sentiment analysis attributes, such as sentiment classification of financial news.
  • FinCUGE (Lu et al., 2023): Involves multiple types of reasoning tasks in the financial field, including numerical analysis and causal relationship extraction.
@@ B.3 The Prompt of Reasoning Selection
For reasoning selection in data filtering step, the prompt template presented in Figure 12 is designed based $7$ key dimensions (Xie et al., 2024b) and is consisted of the evaluation criteria that structure how the reasoning process should be assessed; the inputs, the reasoning process and the standard answer, against which the evaluation is carried out; the description of scoring method, where each criterion is worth one point and the final decision is binary (1 for high-quality reasoning if all seven points are met, otherwise 0), with the score required for standardized extraction.